I was born 3 years before you. Today the sum of our ages is 27. How old are you? How old am I?
Answer by TimothyLamb(4379)

x = your age now
y = my age now
---
y = x + 3
x + y = 27
---
put the system of linear equations into standard form
---
x - y = -3
x + y = 27
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = your age now = 12
y = my age now = 15
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php